What will be conductivity of pure silicon crystal at 300K temperature, if electron hole pairs per cm3 is 1.072x1010 at this temperature, μn = 1350cm2/volt-s and μp = 480 cm2/volt-s?

Grade:12th pass

1 Answers

Sehaj
3 years ago
Conductance=e(nu+n'u') 
=1.6×10^-19×1.07×10^10(1350+480)
=1.6×1.07×10^-9×1830
=3.14×10^-6 mho/cm